# V-HOP: Visuo-Haptic 6D Object Pose Tracking

### v1

@@ 7 Limitation
We follow the model-based object pose tracking setting, which assumes that a CAD model is available for the object. One potential direction to overcome this limitation is to simultaneously reconstruct the object and perform pose tracking, as demonstrated in methods like BundleSDF Wen et al. (2023).
8 Conclusion
@@ Acknowledgments
This work is supported by the National Science Foundation (NSF) under CAREER grant #2143576, grant #2346528, and ONR Grant #N00014-22-1-259. We thank Ying Wang, Tao Lu, Zekun Li, and Xiaoyan Cong for their valuable discussions.
@@ Appendix B Experiments
Figure 12: Accuracy-threshold curve Xiang et al. (2018) on our dataset. V-HOP consistently demonstrates stronger or similar performance as FoundationPose (FP) under various thresholds.

### v2

@@ VII Limitation
We follow the model-based object pose tracking setting, which assumes that a CAD model is available for the object. While assuming a CAD model may limit generalization in in-the-wild applications, it is a well-established assumption in industrial settings, such as warehouses or assembly lines [3, 54]. One potential direction to overcome this limitation is to simultaneously reconstruct the object and perform pose tracking, as demonstrated in methods like BundleSDF [69] and NeuralFeels [54], which offer promising and compatible ways to supply a model to our approach.
VIII Conclusion
@@ Acknowledgments
This work is supported by the National Science Foundation (NSF) under CAREER grant #2143576, grant #2346528, and the Office of Naval Research (ONR) grant #N00014-22-1-259. We thank Ying Wang, Tao Lu, Zekun Li, and Xiaoyan Cong for their valuable discussions. We thank the area chair and the reviewers for providing constructive feedback on improving the quality and clarity of our paper. This research was conducted using computational resources and services at the Center for Computation and Visualization, Brown University.
@@ Appendix B Experiments
Figure 12: Accuracy-threshold curve [72] on our dataset. V-HOP consistently demonstrates stronger or similar performance as FoundationPose (FP) under various thresholds.